# Quality-Aware Decoding: Unifying Quality Estimation and Decoding

### v2

@@ 1 Introduction
In this work, we propose adapting LLM-based MT models to perform QE on partial translations and incorporating this model into the decoding. We create a token-level synthetic QE dataset using WMT Multidimensional Quality Metrics (MQM) data (Burchardt, 2013; Freitag et al., 2024). We then adapt a uni-directional LLM-based MT model to predict whether a token is Good or Bad. Training QE models on these token-level tasks alleviates the data challenge and allows us to exploit the MQM data while simultaneously making the task easier for the model compared to predicting a score directly.
@@
Furthermore, integrating the QE model into NMT during decoding is not trivial, as we need to combine the QE estimates during decoding. Therefore, we use the decoding strategy from Koneru et al. (2024), and modify it to incorporate token-level predictions efficiently with the adapted QE model to provide real-time feedback during the decoding process. We summarize our main findings and contributions below.
@@ 2 Quality-Aware Decoding
The primary objective of this paper is to achieve Quality-Aware Decoding in MT. To accomplish this, it is essential to predict the quality of partial translations and integrate this information during the decoding process. Our approach proposes using one NMT model for generating translations and another adapted NMT model to predict the quality of the candidate translations produced by the first model.
@@ 2.2 Quality Estimation on Partial Sequences
For training this model, we leverage the WMT MQM data containing error annotations in MT outputs. We can treat tokens before an error as Good and those containing inside an error as Bad. Then, we can train in uni-directional manner where each token’s label is predicted using only the preceding context in the hypothesis. This is crucial as we only have the preceding context to estimate the quality for partial hypothesis.
@@ 2.2.3 Fusing Translation and Quality
Model Beams Re-ranking MetricX ($\downarrow$) XCOMET-XXL ($\uparrow$)
English $\rightarrow$ German
Tower 5 _ 2.52 86.93
Tower 25 XCOMET-XL QE 2.37 87.79
Tower 25 Tower QE 2.38 87.40
Tower + Tower QE 5 (25* for Tower QE) _ 2.12 88.95
Tower + Tower QE 5 (25* for Tower QE) XCOMET-XL QE 2.09 89.08
Chinese $\rightarrow$ English
Tower 5 _ 2.42 88.91
Tower 25 XCOMET-XL QE 2.30 89.49
Tower 25 Tower QE 2.32 89.51
Tower + Tower QE 5 (25* for Tower QE) _ 2.26 89.82
Tower + Tower QE 5 (25* for Tower QE) XCOMET-XL QE 2.24 90.00
Table 1: Translation Quality on WMT23 English $\rightarrow$ German Test set. Both XCOMET and MetricX columns use reference for reporting translation quality where as XCOMET-XL QE does not use for re-ranking.
@@ 3 Experimental Setup
Our proposed approach achieves Quality-Aware decoding by combining an NMT model with a token-level QE model, where we adapt the same NMT for QE by adding a classification head. We use the state-of-the-art NMT model, Tower 7B [^3] Unbabel/TowerInstruct-7B-v0.2 (Alves et al., 2024), which provides high-quality translations and has already been exposed to MQM data during instruction-tuning. This ensures that the gains observed in our approach stem from integrating Quality-Aware decoding into the NMT process, rather than introducing new data. Additional details on training and hyper-parameters are provided in Appendix A.1.
@@ 4 Results
We conduct a series of experiments to validate the effectiveness of Quality-Aware decoding and identify the scenarios where it provides the most benefit. First, we evaluate whether our token-level QE model can better estimate sequence quality compared to the log probabilities of the NMT model. Next, we assess the impact of Quality-Aware decoding by comparing it with other approaches to determine if it improves translation quality. We also perform an ablation study to examine whether training the QE model on errors from the same NMT model enhances its performance. Finally, we explore the impact of source sentence length to highlight the limitations of N-best list re-ranking.
@@ 5 Related Work
Integrating QE in NMT: Several advancements have been made in improving QE for NMT over the years (Rei et al., 2021, 2022; Blain et al., 2023; Zerva et al., 2024; Guerreiro et al., 2024). These developments have led to the integration of QE in various ways. One common approach involves applying QE after generating multiple sequences through techniques such as QE re-ranking (Fernandes et al., 2022; Faria et al., 2024) or Minimum Bayes Risk (MBR) decoding (Tomani et al., 2024). Another direction focuses on removing noisy data using QE models, followed by fine-tuning on high-quality data (Xu et al., 2024; Finkelstein et al., 2024). Vernikos and Popescu-Belis (2024) proposes to generate diverse translations as a first step and then combine them. We perform this explicitly by integrating the QE directly into decoding. Recently, Zhang et al. (2024) exploited the MQM data by training models to penalize tokens within an error span, improving translation quality. In contrast, our approach adopts a modular framework, where we propose an expert QE model that is trained independently for targeted training. This modular approach aims to improve performance by decomposing the task into separate translation and QE components.
Reward Modeling in NLG: Quality-Aware decoding shares several similarities with controllable text generation methods, particularly in the use of an additional "Quality/Reward" model that guides the decoding. A well-explored approach for controlling text is altering the decoding with a reward model (Weighted Decoding) (Yang and Klein, 2021). This method modifies the decoding by adjusting token probabilities based on the reward model, allowing for more controlled generation. Similarly, Deng and Raffel (2023) also used a uni-directional reward model, with the aim of maintaining efficiency during generation. This approach minimizes computational complexity while still benefiting from the guiding influence of the reward model. Moreover, recent work by Li et al. (2024) introduced a token-level reinforcement learning-based reward model, providing more fine-grained feedback that enhances control over text generation at a granular level. While similar, the key contribution in our work lies in the development of the first uni-directional QE model for translation.
@@ 6 Conclusion
We have shown the importance of Quality-Aware decoding to improve translation quality, rather than relying on post-generation techniques. In this work, we demonstrated how MQM data can be used to build a uni-directional token-level QE model, which is then integrated into the decoding process. Through a series of experiments, we showed that our Quality-Aware decoding approach results in measurable improvements in translation quality. Notably, we did not introduce new training data to the NMT model, and show that the gains stem from Quality-Aware decoding.
@@ 7 Limitations
While our Quality-Aware decoding improves translation quality, it adds considerable computational complexity to the inference process. Theoretically, this approach would double the time needed to generate a translation and require additional memory to utilize the token-level QE model. One potential solution to mitigate this issue could be to use token-level QE as a reward model for training via Reinforcement Learning.
Additionally, we trained our model on a limited set of human-annotated MQM data. However, current QE models, such as XCOMET, are capable of predicting error tags using the reference with reasonable quality. This suggests that further improvements could be achieved if these models were trained on larger-scale datasets, providing more nuanced feedback and refining translation quality even further.
@@ A.2 Partial vs Full Sequence Quality Estimation
We see that the both Tower QE models based on full sentences outperforms the partial model. However, this is expected as it has seen the entire context and was also trained on larger amounts of data. Nonetheless, the partial model still achieves much higher correlaiton that the log probabilities showcasing its potential for Quality-Aware decoding.
@@ A.3 Robustness to re-ranking weight
In our method, we introduce a hyperparameter, $\alpha$, to merge the probabilities from the token-level QE model and the translation model. This section analyzes the impact of $\alpha$ on the final translation quality.
To efficiently evaluate its effect, we re-rank the N-best list using different values of $\alpha$. This approach allows us to estimate the ideal value of $\alpha$ without the need for joint decoding multiple times. If the re-ranking model (in this case, Tower QE) is beneficial, we expect that any $\alpha$ less than 1 will improve translation quality, as it demonstrates that incorporating the probabilities from the QE model is helpful.
We visualize this impact in Figure 4. The results show that using an $\alpha$ less than 1 leads to improved translation quality in both scenarios. This indicates that relying entirely on the NMT model does not yield the best results and highlights the importance of the Tower QE model.
Thus, we emphasize that re-ranking the N-best list provides an effective way to tune the value of $\alpha$, and it remains robust to different values.
Tower Token-Level QE Prompt English:{src_sent} German: {tgt_sent}
(a) English $\rightarrow$ German
(b) Chinese $\rightarrow$ English

### v3

@@ 1 Introduction
In this work, we propose adapting LLM-based NMT models to perform QE on partial translations and incorporating this model into the decoding. We create a token-level synthetic QE dataset using WMT Multidimensional Quality Metrics (MQM) data (Burchardt, 2013; Freitag et al., 2024). We then adapt a uni-directional LLM-based MT model to predict whether a token is GOOD or BAD. Training QE models on these token-level tasks alleviates the data challenge and allows us to exploit the MQM data while simultaneously making the task easier for the model compared to predicting a score directly.
@@
Furthermore, integrating the QE model into NMT during decoding is not trivial, as we need to combine the QE estimates during decoding. Therefore, we modify the decoding strategy from Koneru et al. (2024) to incorporate token-level predictions efficiently with the adapted QE model to provide real-time feedback during the decoding process. We summarize our contributions below.
@@ 2 Quality-Aware Decoding
The primary objective of this paper is to achieve Quality-Aware Decoding in NMT. To accomplish this, it is essential to predict the quality of partial translations and integrate this information during the decoding process. Our approach proposes using one NMT model for generating translations and another adapted NMT model to predict the quality of the candidate translations produced by the first model.
@@ 2.2 Quality Estimation on Partial Sequences
For training this model, we leverage the WMT MQM data containing error annotations in NMT outputs. We can treat tokens before an error as GOOD and those containing inside an error as BAD. Then, we can train in uni-directional manner where each token’s label is predicted using only the preceding context in the hypothesis. This is crucial as we only have the preceding context to estimate the quality for partial hypothesis.
@@ 2.2.3 Fusing Translation and Quality
Model Beams Re-ranking MetricX ($\downarrow$) XCOMET-XXL ($\uparrow$)
English $\rightarrow$ German
Tower 5 _ 2.52 86.93
Tower 25 XCOMET-XL QE 2.37 87.79
Tower 25 Tower QE 2.38 87.40
Tower + Tower QE 5 (25* for Tower QE) _ 2.12 88.95*
Tower + Tower QE 5 (25* for Tower QE) XCOMET-XL QE 2.09* 89.08*
Chinese $\rightarrow$ English
Tower 5 _ 2.42 88.91
Tower 25 XCOMET-XL QE 2.30 89.49
Tower 25 Tower QE 2.32 89.51
Tower + Tower QE 5 (25* for Tower QE) _ 2.26 89.82*
Tower + Tower QE 5 (25* for Tower QE) XCOMET-XL QE 2.24* 90.00*
Table 2: Translation Quality on WMT23 English $\rightarrow$ German Test set. Both XCOMET and MetricX columns use reference for reporting translation quality where as XCOMET-XL QE does not use for re-ranking. Best scores according to each metric are highlighted in bold. We report the top cluster indicated with asterisks and are no worse than other systems determined by Paired T-Test and bootstrap resampling with p < 0.05
@@ 3 Experimental Setup
Our proposed approach achieves Quality-Aware decoding by combining an NMT model with a token-level QE model, where we adapt the same NMT for QE by adding a classification head. We use the state-of-the-art NMT model, Tower 7B [^3] Unbabel/TowerInstruct-7B-v0.2 (Alves et al., 2024), which provides high-quality translations and has already been exposed to MQM data during instruction-tuning. This ensures that the gains observed in our approach stem from integrating Quality-Aware decoding into the NMT process, rather than introducing new data. We find $\alpha$ by setting it to the optimal re-ranking weight on the validation set ( See Appendix A.3 for details). Additional details on training the QE models and hyper-parameters during decoding are provided in Appendix A.1.
@@ 4 Results
We conduct a series of experiments to validate the effectiveness of Quality-Aware decoding and identify the scenarios where it provides the most benefit. First, we evaluate whether our token-level QE model can better estimate sequence quality compared to the log probabilities of the NMT model. Next, we assess the impact of Quality-Aware decoding by comparing it with other approaches to determine if it improves translation quality. We also perform an ablation study to examine whether training the QE model on errors from the same NMT model enhances its performance. Then, we explore the impact of source sentence length to highlight the limitations of N-best list re-ranking. Finally, we compare our proposed approach with existing Quality-Aware decoding strategies and also their inference time to highlight the latency/quality trade-off.
@@ 4.4 Sentence vs Document-level Translation
Model Beams
Re-ranking/
Utility Metric
MetricX XCOMET-XXL
Tower 25 XCOMET-XL 2.37 87.79
Tower + MBR 25 wmt22-comet-da 2.75 87.53
Tower + QE-Fusion 25 XCOMET-XL 2.38 87.76
Tower + Tower QE (Ours) 5(25) _ 2.12 88.95*
Tower + Tower QE (Ours) 5(25) XCOMET-XL 2.09* 89.08*
Tower + Tower QE (Ours) + QE-Fusion 5(25) XCOMET-XL 2.10 89.03*
Table 5: Translation quality on WMT 23 English $\rightarrow$ German. We report the top cluster indicated with asterisks and are no worse than other systems determined by Paired T-Test and bootstrap resampling with p < 0.05. For MBR, we use wmt22-comet-da as the utility metric wheras XCOMET-XL for QE-Fusion (Vernikos and Popescu-Belis, 2024).
Model Beams
Avg Time
XCOMET-XXL
Tower 5 5.20 86.93
Tower
XCOMET-XL Rerank
25 13.16 + 1.78 87.79
Tower
MBR wmt22-comet-da
25 13.16 + 0.34 87.56
Tower
Tower QE (Ours)
5(25) 21.04 88.94
Table 6: Average inference time on WMT 23 English-German document-level test set. For MBR and re-ranking, $13.16$ is the time used for generating 25 candidates.
4.5 Compatibility with Sampling-based Strategies
Our proposed approach integrates the feedback during decoding time to generate high quality N-best list. Therefore, it can be further combined with strategies such as Minimum Bayes Risk (MBR) (Freitag et al., 2022) or QE-Fusion (Vernikos and Popescu-Belis, 2024) decoding that only rely on the sampled candidates. We compare the different decoding strategies and report the scores in Table 5. For MBR decoding, we do epsilon sampling (epsilon=0.02) and generate 25 candidates with wmt22-comet-da as utility metric. We do not sample more as it is expensive especially at document-level with 7B model and requires multiple runs. For QE fusion, we use XCOMET-XL as the utility metric. We perform this on English-German at paragraph-level as our hypothesis is that the N-best list is problematic for long sequences.
We find that decoding with Tower QE is significantly better than the other approaches in this setting according to XCOMET-XXL metric. We also would like to highlight that it is compatible with sampling based approaches and show that by performing QE-fusion on the top 5 beams with our decoding approach. While the scores are slightly lower, human evaluation is necessary to compare these systems.
4.6 Inference Time
While the quality of translation improves, our approach also is computationally more expensive. To demonstrate this, we compare different decoding strategies and report the latency and quality in Table 6. To calculate the latency, we take the average time in seconds for inference on the WMT 23 English $\rightarrow$ German Test set.
Note that MBR with XCOMET-XL is extremely expensive given the large size of the QE model and the amount of long samples (25*24*557 examples at paragraph level limit batching) that need to be passed through the model due to the exponential nature of MBR. We see that due to the uni-directional nature of the Tower QE, the total inference time is less than double. Further, re-ranking with XCOMET-XL is the fastest given that a single forward pass is needed after generation.
@@ 5 Related Work
Integrating QE in NMT: Several advancements have been made in improving QE for NMT over the years (Rei et al., 2021, 2022; Blain et al., 2023; Zerva et al., 2024; Guerreiro et al., 2024). These developments have led to the integration of QE in various ways. One common approach involves applying QE after generating multiple sequences through techniques such as QE re-ranking (Fernandes et al., 2022; Faria et al., 2024) or Minimum Bayes Risk (MBR) decoding (Tomani et al., 2024). Another direction focuses on removing noisy data using QE models, followed by fine-tuning on high-quality data (Xu et al., 2024; Finkelstein et al., 2024). Vernikos and Popescu-Belis (2024) proposes to generate diverse translations as a first step and then combine them. We perform this explicitly by integrating the QE directly into decoding. Recently, Zhang et al. (2024) exploited the MQM data by training models to penalize tokens within an error span, improving quality. In contrast, our approach adopts a modular framework, where we propose an expert QE model that is trained independently for targeted training. This modular approach aims to improve performance by decomposing the task into separate translation and QE components.
Reward Modeling in NLG: Quality-Aware decoding shares similarities with controllable text generation, particularly in using a "Quality/Reward" model to guide decoding. Methods like Weighted Decoding (Yang and Klein, 2021) adjust token probabilities for controlled generation, while Deng and Raffel (2023) use a uni-directional reward model to maintain efficiency. Li et al. (2024) further enhance control with a token-level reinforcement learning-based model. While related, our key contribution is the development of the first uni-directional QE model specifically for translation.
@@ 6 Conclusion
We demonstrated the value of Quality-Aware decoding in improving translation quality without relying on post-generation methods. Using MQM data, we built a uni-directional token-level QE model and integrated it into the decoding process. Our experiments show measurable quality gains, achieved without adding new training data to the NMT model, highlighting that improvements come only from the decoding approach.
@@ 7 Limitations
Although our Quality-Aware decoding improves translation quality, it adds considerable computational complexity to the inference process. Theoretically, this approach would double the time required to generate a translation and would require additional memory to utilize the token-level QE model. One potential solution to mitigate this issue could be to use token-level QE as a reward model for training through reinforcement learning.
Furthermore, we trained our model on a limited set of human-annotated MQM data. However, current QE models, such as XCOMET, are capable of predicting error tags using the reference with reasonable quality. This suggests that further improvements could be achieved if these models were trained on larger-scale datasets, providing more nuanced feedback and refining translation quality even further.
In addition, human evaluation is necessary to validate if the translation quality also improves with human judgment. Although we were able to better integrate MQM data during decoding, it decreases confidence in relying completely on automatic metrics.
@@
Acknowledgments
Part of this work was performed on the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Württemberg and by the Federal Ministry of Education and Research.
@@ A.2 Partial vs Full Sequence Quality Estimation
We see that the both Tower QE models based on full sentences outperforms the partial model. However, this is expected as it has seen the entire context and trained on larger amounts of data. Still, the partial model achieves much higher correlaiton that the log probabilities showcasing its potential for Quality-Aware decoding.
@@ A.3 Robustness to re-ranking weight
We introduce a hyperparameter, $\alpha$, to merge probabilities from the token-level QE model and the translation model. To analyze its impact, we re-rank the N-best list using various $\alpha$ values, avoiding repeated joint decoding. If the QE model is helpful, we expect translation quality to improve when $\alpha<1$.
Figure 4 shows that lower $\alpha$ values consistently yield better results, confirming that incorporating QE probabilities improves translation. This highlights the value of Tower QE and shows that re-ranking is an effective and robust way to tune $\alpha$.
@@
Tower Token-Level QE Prompt English:{src_sent} German: {tgt_sent}
(a) English $\rightarrow$ German
(b) Chinese $\rightarrow$ English